1) A rock is thrown downwards with a speed of 10 m/o from the top of a bridge 20m above water (a) Determine the time it takes to reach the water (b) Determine its velocity as it hits the water.

I need help asap
Answer #1

here,

1)

the initial speed , u = 10 m/s

a)

let the time taken be t

h = 0 + 0.5 * g * t^2

20 = 0 + 0.5 * 9.81 * t^2

solving for t

t = 2.02 s

the time taken is 2.02 s

b)

the velocity of rock when it hits the water , v = u + g * t

v = 10 + 9.81 * 2.02 m/s

v = 29.8 m/s